# Causal Mechanisms of Subpolar Gyre Variability in CMIP6 Models

### v2

@@ Abstract
The subpolar gyre is at risk of crossing a tipping point under future climate change associated with the collapse of deep convection. As such tipping can have significant climate impacts, it is important to understand the mechanisms at play and how they are represented in modern climate models. In this study we use causal inference to investigate the representation of several proposed mechanisms of subpolar gyre variability in CMIP6 models. As expected, an increase in sea surface salinity or a decrease in sea surface temperature leads to an increase in mixed layer depth in nearly all CMIP6 models due to an intensification of deep convection. However, the effect of convection to modify sea surface temperature due to restratification is less well captured. In most models the deepening of the mixed layer caused by an increase of sea surface salinity, does result in a cooling of the water at intermediate depths. The feedback from the subsurface temperature through density to the strength of the subpolar gyre circulation is more ambiguous, with fewer models indicating a significant link. Those that do show a significant link, do not agree on its sign. One model (CESM2) contains all proposed mechanisms, with both a negative and delayed positive feedback loop.
@@ 1 Introduction
In CMIP6 models, deep convection is driven by vertical density differences and occurs when the surface density exceeds that of the waters below. Density is controlled by both temperature and salinity, where either an increase in sea surface salinity (SSS) or a decrease in sea surface temperature (SST) can lead to convection, deepening the mixed layer depth (MLD). These two causes can be represented as in Figure 1 by the orange and brown (A1) arrows in the yellow box. It is still a topic of debate which of the two (SST or SSS) is more important for driving variability in convection, and the results differ depending on the timescales of interest. Some studies indicate SSS as the dominant driver of convection (e.g. Hátún et al., 2005; Gelderloos et al., 2012; Yamamoto et al., 2020), where changes in salinity are primarily linked to ocean transport. Others show the importance of surface heat fluxes, which affect SST, on the strength of convection (e.g. Yashayaev and Loder, 2016; Piron et al., 2017). On interannual timescales, SSS is a stronger driver than SST (e.g. Yamamoto et al., 2020).
Convection also feeds back to the surface as water from depth is mixed with the surface water, lowering the surface density. The effect of this re-stratification on temperature has been found to be significant, whereas the effect on salinity is not as strong (Lazier et al., 2002). This can lead to bistability of convective and non-convective states as shown in conceptual models (Welander, 1982) and is represented by the orange and pink (A2) arrows in Figure 1. Even without detailed marginal sea processes representations in CMIP6 models, the density in the SPG region affects the strength of the (horizontal) gyre circulation, which in turn alters SST and SSS. This way changes in convection occurring can alter the strength of the SPG circulation, providing a possible feedback loop Born and Mignot (2012).
@@ 2.1 Data
To test for the presence of a mechanism in CMIP6 data, it is desirable that as little confounding or forcing mechanisms are present. Therefore, the piControl runs of CMIP6 are considered (Eyring et al., 2016). The variables of interest in the CMIP6 database are the sea surface salinity (sos), sea surface temperature (tos), mixed layer depth (mlotst), subsurface potential temperature (thetao) and the barotropic streamfunction (msftbarot). The sea surface temperature and sea surface salinity are used to compute the density following the TEOS-10 equation of state (Roquet et al., 2015). The barotropic streamfunction is used as a measure of the strength of the subpolar gyre because it is hard to determine the baroclinic part of the flow and the use of density variables to this end would build in a link between density and the gyre strength. With these five variables we compute (scalar) time series for all six feedback elements shown in Figure 1. The regions used for this computation are detailed in the following paragraphs. From hereon we use SPG for the strength of the subpolar gyre, SSS for the sea surface salinity, SST for the sea surface temperature, MLD for the mixed layer depth, SubT for the temperature at intermediate depth and Rho for the density. For the mechanisms (A) linked to convection we consider all 47 CMIP6 models for which at least 100 years of data is available (list in Figure 2). Not all of these models have the barotropic streamfunction available as a variable and thus we are left with 32 models for the analysis of mechanism (B) (see SI for the list).
@@
Because of the colder temperatures, and resulting lower surface density, convection primarily occurs in winter (Birol Kara et al., 2003; Heuzé, 2017). Therefore, we focus our analysis on the winter months and compute the mean of each variable over January-February-March. Convection is a local process and the location where it is strongest differs between CMIP6 models, as can be seen in Figure 2. For the initial analysis on the mechanism of convection (A) we therefore select a different region for each model to compute the SSS, SST and MLD indices. For each model a 5∘ by 5∘ box around the location of maximum winter (JFM) mixed layer depth is obtained (restricted to the SPG region as indicated by the dashed line), and the indices are computed as the spatial average over those boxes. This ensures that for each model we focus on the region where that model has the strongest convection.
@@
For the analysis of the Born hypothesis (B) one of the relevant variables is the barotropic streamfunction, as an indicator of the strength of the subpolar gyre. The interaction between the strength of the circulation and the temperature and salinity of the water differs between locations, since the water flowing into the considered region has a different source and thus different properties. This makes it difficult to compare models when using a separate box for each model. Therefore, we decide to work with one fixed box for all models when studying this mechanism. The region studied is chosen based on where convection takes place in CMIP6 models. For this we consider MLD as an indicator, assuming deep convection takes place if it exceeds 1000m (Marshall and Schott, 1999). In Figure 2 the number of models in which deep convection takes place during at least 100 model years is shown. It indicates that the Labrador sea is the region in the subpolar gyre where deep convection takes place in most models. Therefore, the region considered in this study is 54-63∘N, 47-60∘W, indicated by the solid box in Figure 2. The spatial average over this box is computed for all variables considered. For SubT we consider values between 50m and 1000m depth (as in Born and Mignot (2012)).
@@ 2.2 Causal Links and Causal Effect
To verify whether the convection and Born mechanisms are represented in CMIP6 models we analyse the interaction between variables in two ways for each model. Firstly, causal inference is used to identify whether a connection between two variables is significant (at the 5%-level). Secondly, causal effect determines the strength of each link given a network of links. Both approaches are conducted using the Tigramite package for python. For the identification of significant links, the Peter and Clark momentary conditional independence (PCMCI) algorithm is used (Runge et al., 2014, 2019b, 2019a). This is a causal discovery algorithm which identifies significant causal links between the input variables for a given set of lag-times. Several versions of the algorithm have been developed, of which we use PCMCI+ because it can give direction to contemporaneous (no lag) links (Runge, 2020).
The PCMCI algorithm allows to infer causal links between variables from time series data. Following Pearl et al. (2016) a variable causes if is non-zero, where is called the parent, the child and is an intervention. When considering time-series data it is not possible to intervene, i.e. , and therefore a time-lag is used to infer causality, assuming the cause happens prior to the effect. This estimation of the links by the algorithm requires a number of assumptions to be valid, of which a full description is given in Runge (2018). Here we briefly discuss the ones that are most relevant to this study.
Firstly, causal sufficiency is assumed. This means that all causally relevant variables are included and there exist no other (unobserved) variables that influence a pair of variables in the considered set. Secondly, the time series is assumed to be stationary, in the sense that the presence and strength of the links does not change in time. The last assumption we mention here is linearity of the links, which strictly speaking is unlikely to be valid. There are multiple metrics that can be used in the algorithm, but the more non-linear they get, the longer it takes too compute, making them unfeasible for our application to many models. Furthermore, linearity has been shown to be a good first order approximation for climate data (Kretschmer et al., 2021; Di Capua et al., 2020). Therefore, we use partial correlation as our metric, extending the commonly used frame of correlation by taking into account the effect of auto-correlation and (included) confounding factors (Saggioro and Shepherd, 2019). To be specific, we use a robust partial correlation measure (RobustParCorr) which is more suitable for non-Gaussian data. This choice is made because especially the mixed layer depth does not follow a Gaussian distribution in most models.
@@ 3 Mechanisms of Convection
We start by studying the two convection mechanisms (A) represented in Figure 1. Because of the close relationship between SSS and SST, causal inference struggles to identify the expected links, especially for direct interaction, and shows a large spread between models (see SI). The absence of common drivers, such as the circulation strength of the gyre, among the included variables means that the assumption of causal sufficiency is violated, which makes it impossible to robustly identify the presence of links in the models. Therefore, we limit our analysis to the computation of the causal effects between the variables. The causal effects between the variables were obtained by using the links for mechanisms (A1) and (A2) as shown in Figure 1 and assuming a lag-1 memory effect of each of the variables onto itself. The latter is a feature clearly identified in most models using causal inference. This part of the analysis may be hindered by the violation of causal sufficiency as well, but it can still provide valuable insight in the interaction mechanisms.
@@
The results for this analysis of causal effects are shown in Figure 3. For mechanism (A1) we find that indeed an increase in SSS leads to more convection, i.e. an increase in MLD. A decrease in SST has the same effect, although slightly weaker in most models. The results for mechanism (A2) are less pronounced. In most models an increase in SSS causes an increase in MLD, although the effect is weaker than in (A1). The feedback from MLD to SST is unclear, with some models indicating a positive effect, in line with theory, but others showing a negative causal effect. It shows that the effect that convection has on SST is hard to detect. This is likely due to confounding factors (violating causal sufficiency), such as the ocean circulation, that also strongly impact SST (and SSS) and that way obscure the signal. The reason this affects hypothesis (A2) and not (A1), is that in (A2) SST is a child of SSS (through MLD). In (A1) such a relation is absent and thus confounding factors, i.e. parents of both SST and SSS, do not impact the results. Thus, in order to be able to identify a robust and reliable signal we require the hypothesis to be causally sufficient, by which we mean a hypothesis where all relevant variables are included. One hypothesis for SPG variability that is causally sufficient is that of the mechanism (B) shown in the large loop in Figure 1. It includes all variables relevant to bistable variability of the SPG, and thus can be tested using causal inference. This does not mean that the mechanism in reality is not affected by confounding factors, but the hypothesized mechanism is not and thus can be tested.
@@ 4 Mechanism of Subpolar Gyre Variability
The verification in CMIP6 models of the hypothesized mechanism (B) as shown in Figure 1 is done in two steps. Firstly, causal inference is used to find which of the links are significant, which is done by studying one link at a time. This means we apply causal discovery to , and the causing variable of (), accounting for its possible impact on auto-correlation, to verify the significance of the links. It is unlikely that all models agree on which links in the theoretical model are significant at which lags and thus we start with a discussion of the number of models in which the links are found to be significant in Section 4.1. Based on these results we move to the second step; computation of the causal effect. For this we set the lags in the network as found in Section 4.1 and use this to test the strength of the links (Section 4.2). Note that the computations here are done for the variables considered in the Labrador sea box as shown in Figure 2, which is different from the model dependent areas considered before. Comparison between the two regions shows no significant differences in the causal effects for the convection mechanisms (see SI), indicating the mechanism (A1) can also be identified when considering the same Labrador Sea region for all models.
@@ 4.1 Significant Links
The number of models for which each of the five links is significant are shown in Figure 4 for lags up to 10 years (individual model results are shown in the SI). Results for both the theoretical links (bottom row), as well as the influence of variables on themselves (top row), i.e. their memory, are shown. For contemporaneous links also the sum of the correctly directed and unknown links is shown. Looking at the top row we see that SPG has a long memory in many models, with significant links of up to 5 years lag (lag-5) in more than 5 models. SSS, MLD and Rho have substantially shorter memory, whereas SubT-memory is two years in most models.
For the links of the proposed subpolar gyre variability mechanism (B), we find that the SPGSSS link is absent in 22 out of the 32 models, especially when direct causation is concerned (only three models, one of which also has a link at lag-1). At lag-1 eight models show a significant link. The SSSMLD link is correctly represented in 23 models, either direct or at lag-1 (or both), with 13 models indicating an unknown direction for the contemporaneous link. Here, 13 models show a link in the opposite direction (some with a lag-1 link in the right direction), indicating these two variables are strongly connected in models. The direct MLDSubT link is also found in many models; 22 in this case with half being unoriented. At lag-1, seven models find a significant link. The SubTRho link is not present in many models, with only seven models showing a correctly directed or unoriented contemporaneous link and six models indicating a significant link at lag-1. Lastly, the RhoSPG link is found in seven more models than the SPGSSS link, with eleven models (out of the 17) indicating lag-1 to be significant. In total there are six models that have both links with SPG as significant.
All but one model capture at least one step of the process (B) (only CanESM-1 does not). The best captured links are SSSMLD and MLDSubT, which are found in 23 and 27 CMIP6 models respectively. The links to and from SPG are found in only 10 and 17 models respectively. This is to be expected as this connection relies on the modelling of advection which is dependent on the used parametrisation (Griffies, Stephen, 2018; Small et al., 2020). The SPGSSS link at longer lag is not found. There is one model, CESM2, that captures all links (except the lagged one), which will be discussed in more detail in the following section.

### v3

@@ Abstract
The subpolar gyre is at risk of crossing a tipping point under future climate change associated with the collapse of deep convection. As such tipping can have significant climate impacts, it is important to understand the mechanisms at play and how they are represented in modern climate models. In this study we use causal inference to investigate the representation of several proposed mechanisms of subpolar gyre variability in CMIP6 models. As expected, an increase in sea surface salinity or a decrease in sea surface temperature leads to an increase in mixed layer depth in nearly all CMIP6 models due to an intensification of deep convection. However, the effect of convection to modify sea surface temperature due to re-stratification is less clear. In most models the deepening of the mixed layer caused by an increase of sea surface salinity, does result in a cooling of the water at intermediate depths. The feedback from the subsurface temperature through density to the strength of the subpolar gyre circulation is more ambiguous, with fewer models indicating a significant link. Those that do show a significant link, do not agree on its sign. The CMIP6 models that have the expected sign for the links from density to the subpolar gyre strength and on to sea surface salinity, are also the models in which abrupt shifts in the subpolar gyre region have been found in climate change scenario runs. One model (CESM2) contains all proposed mechanisms, with both a negative and delayed positive feedback loop being significant.
@@ 1 Introduction
In CMIP6 models, deep convection is driven locally by vertical density differences and occurs when the surface density exceeds that of the waters below. Density is controlled by both temperature and salinity, where either an increase in sea surface salinity (SSS) or a decrease in sea surface temperature (SST) can lead to convection, deepening the mixed layer depth (MLD). These two causes can be represented as in Figure 1 by the orange and brown (A1) arrows in the yellow box. It is still a topic of debate which of the two (SST or SSS) is more important for driving variability in convection, and the results differ depending on the timescales of interest. Some studies indicate SSS as the dominant driver of convection (e.g. Hátún et al., 2005; Gelderloos et al., 2012; Yamamoto et al., 2020), where changes in salinity are primarily linked to ocean transport. Others show the importance of surface heat fluxes, which affect SST, on the strength of convection (e.g. Yashayaev and Loder, 2016; Piron et al., 2017). On interannual timescales, SSS is a stronger driver than SST (e.g. Yamamoto et al., 2020).
Convection also feeds back to the surface as warmer water from depth is mixed with the surface water, lowering the surface density. The effect of this re-stratification on temperature has been found to be significant, whereas the effect on salinity is not as strong (Lazier et al., 2002). This can lead to bistability of convective and non-convective states as shown in conceptual models (Welander, 1982) and is represented by the orange and pink (A2) arrows in Figure 1. Even without detailed marginal sea processes representations in CMIP6 models, the density in the SPG region affects the strength of the (horizontal) gyre circulation, which in turn alters SST and SSS. This way changes in convection occurring can alter the strength of the SPG circulation, providing a possible feedback loop (Born and Mignot, 2012).
@@ 2.1 Data
To test for the presence of a mechanism in CMIP6 data, it is desirable that as little confounding or forcing mechanisms are present. Therefore, the piControl runs of CMIP6 are considered (Eyring et al., 2016). The variables of interest in the CMIP6 database are the sea surface salinity (sos), sea surface temperature (tos), mixed layer depth (mlotst), subsurface potential temperature (thetao) and the barotropic streamfunction (msftbarot). The sea surface temperature and sea surface salinity are used to compute the density following the TEOS-10 equation of state (Roquet et al., 2015). The barotropic streamfunction is used as a measure of the strength of the subpolar gyre because it is hard to determine the baroclinic part of the flow and the use of density variables to this end would build in a link between density and the gyre strength. With these five variables we compute (scalar) time series for all six feedback elements shown in Figure 1. The regions used for this computation are detailed in the following paragraphs. For the temperature at intermediate depth we take into account depth levels between 50m and 1000m (as in Born and Mignot (2012)). From hereon we use SPG for the strength of the subpolar gyre, SSS for the sea surface salinity, SST for the sea surface temperature, MLD for the mixed layer depth, SubT for the temperature at intermediate depth and Rho for the density. For the mechanisms (A) linked to convection we consider all 47 CMIP6 models for which at least 100 years of data is available (list in Figure 2). Not all of these models have the barotropic streamfunction available as a variable and thus we are left with 32 models for the analysis of mechanism (B) (see Tab. S1 for the list).
@@
Because of the colder temperatures, and resulting lower surface density, convection primarily occurs in winter (Birol Kara et al., 2003; Heuzé, 2017). Therefore, we focus our analysis on the winter months and compute the mean of each variable over January-February-March. Convection is a local process and the location where it is strongest differs between CMIP6 models, as can be seen in Figure 2. For the initial analysis on the mechanisms of convection (A) we therefore select a different region for each model to compute the SSS, SST and MLD indices. For each model a 5∘ by 5∘ box around the location of maximum winter (JFM) mixed layer depth is obtained (restricted to the SPG region as indicated by the dashed line), and the indices are computed as the spatial average over those boxes. This ensures that for each model we focus on the region where that model has the strongest convection.
@@
For the analysis of the Born hypothesis (B) one of the relevant variables is the barotropic streamfunction, as an indicator of the strength of the subpolar gyre. The interaction between the strength of the circulation and the temperature and salinity of the water differs between locations, since the water flowing into the considered region has a different source and thus different properties. This makes it difficult to compare models when using a separate local box for each model. Therefore, we decide to work with one fixed box for all models when studying this mechanism. The region studied is chosen based on where convection takes place in CMIP6 models. For this we consider MLD as an indicator, assuming deep convection takes place if it exceeds 1000m (Marshall and Schott, 1999). In Figure 2 the number of models in which deep convection takes place during at least 100 model years is shown. It indicates that the Labrador sea is the region in the subpolar gyre where deep convection takes place in most models. Therefore, the region considered in this study is 54-63∘N, 47-60∘W, indicated by the solid box in Figure 2. The spatial average over this box is computed for all variables considered. Despite not all models showing (deep) convection in this region, this approach does allow for a better comparison of mechanism (B) between models because of the inclusion of the strength of the subpolar gyre. Note that this means that the convection mechanisms (A) are considered local for each model, while mechanism (B) is studied in the whole Labrador Sea area.
@@ 2.2 Causal Links and Causal Effect
To verify whether the convection and Born mechanisms are represented in CMIP6 models we analyse the interaction between variables in two ways for each model. Firstly, causal inference is used to identify whether a connection between two variables is significant (at the 5%-level). Secondly, causal effect determines the strength of each link given a network of links. Both approaches are conducted using the Tigramite package for Python. For the identification of significant links, the Peter and Clark momentary conditional independence (PCMCI) algorithm is used (Runge et al., 2014, 2019b, 2019a). This is a causal discovery algorithm which identifies significant causal links between the input variables for a given set of lag-times. Several versions of the algorithm have been developed, of which we use PCMCI+ because it can give direction to instantaneous (no lag) links (Runge, 2020).
The PCMCI algorithm allows to infer causal links between variables from time series data. Following Pearl et al. (2016) a variable causes if is non-zero, where is called the parent, the child and is an intervention. In an experimental setting such as a wave tank, an intervention could be to change the frequency of the forcing to study the effect on wave height. However, in reality such interventions, i.e. , are not possible and in high-resolution climate models they are computationally too expensive and infeasible. Thus, for the time-series data that is available, a time-lag is used to infer causality, assuming the cause happens prior to the effect. This estimation of the links by the algorithm requires a number of assumptions to be valid, of which a full description is given in Runge (2018). Here we briefly discuss the ones that are most relevant to this study.
Firstly, causal sufficiency is assumed. This means that all causally relevant variables are included and there exist no other (unobserved) variables that influence a pair of variables in the considered set. Secondly, the set of causal links is assumed to be stationary, in the sense that the presence and strength of the links does not change in time. The last assumption we mention here is linearity of the links, which strictly speaking is unlikely to be valid. There are multiple metrics that can be used in the algorithm, but the more non-linear they get, the longer it takes to compute, making them unfeasible for our application to many models. Furthermore, linearity has been shown to be a good first order approximation for climate data (Kretschmer et al., 2021; Di Capua et al., 2020). Therefore, we use partial correlation as our metric, extending the commonly used frame of correlation by taking into account the effect of auto-correlation and (included) confounding factors (Saggioro and Shepherd, 2019). To be specific, we use a robust partial correlation measure (RobustParCorr) which is more suitable for non-Gaussian data. This choice is made because especially the mixed layer depth does not follow a Gaussian distribution in most models.
@@ 3 Mechanisms of Convection
We start by studying the two (local) convection mechanisms (A) represented in Figure 1. Because of the close relationship between SSS and SST, causal inference struggles to identify the expected links, especially for instantaneous interaction, where there is a large spread between models (see Fig. S1). For example, in many models a causal link between SST and SSS is identified due to the absence of common drivers, such as the circulation strength of the gyre. Not including these possible drivers means that the assumption of causal sufficiency is violated, making it impossible to robustly identify the presence of links in the models. Therefore, we limit our analysis to the computation of the causal effects between the variables. The causal effects between the variables are obtained by using the links for mechanisms (A1) and (A2) as shown in Figure 1. We consider both links in (A1) to be instantaneous, while for (A2) we take the SSSMLD link at lag-0 and the MLDSST link at lag-1 because of the slower effect of re-stratification (see also Fig. S1). Furthermore, we assume a lag-1 memory effect of each of the variables onto itself, which is a feature clearly identified in most models using causal inference. This part of the analysis may be hindered by the violation of causal sufficiency as well, but it can still provide valuable insight into the local interaction mechanisms.
@@
The results for this analysis of causal effects are shown in Figure 3. For mechanism (A1) we find that indeed an increase in SSS leads to more convection, i.e. an increase in MLD. A decrease in SST has the same effect, although slightly weaker in most models. The results for mechanism (A2) are less pronounced. In most models an increase in SSS causes an increase in MLD, although the effect is weaker than in (A1). The difference with (A1) arises because in the computation of the causal effect one conditions on all parents of MLD; in the case of (A1) those are SST and MLD at lag-1, whereas in the case of (A2) it is only MLD at lag-1. The identified feedback from MLD to SST is not in line with the theory, with most models indicating a negative causal effect. This is likely due to confounding factors (violating causal sufficiency), such as the ocean circulation, that also strongly impact SST (and SSS) and that way obscure the signal. The reason this affects hypothesis (A2) and not (A1), is that in (A2) SST is a child of SSS (through MLD). In (A1) such a relation is absent and thus confounding factors, i.e. parents of both SST and SSS, do not impact the results. Thus, in order to be able to identify a robust and reliable signal we require the hypothesis to be causally sufficient, by which we mean a hypothesis where all relevant variables are included. One hypothesis for SPG variability that is causally sufficient is that of the mechanism (B) shown in the large loop in Figure 1. It includes all variables relevant to bistable variability of the SPG, and thus can be tested using causal inference. This does not mean that the mechanism in reality is not affected by confounding factors, but the hypothesized mechanism is not and thus can be tested. Adding for example only the strength of the subpolar gyre to the local analysis of convection processes discussed here, would lead to results that are hard to interpret since the location of deepest convection differs strongly between models (as discussed in Section 2.1).
@@ 4 Mechanism of Subpolar Gyre Variability
The verification in CMIP6 models of the hypothesized mechanism (B) as shown in Figure 1 is done in two steps. Firstly, causal inference is used to find which of the links are significant, which is done by studying one link at a time. This means we apply causal discovery to , and the causing variable of (), accounting for its possible impact on auto-correlation, to verify the significance of the links. It is unlikely that all models agree on which links in the theoretical model are significant at which lags and thus we start with a discussion of the number of models in which the links are found to be significant in Section 4.1. Based on these results we move to the second step; computation of the causal effect. For this we set the lags in the network as found in Section 4.1 and use this to test the strength of the links (Section 4.2). Note that the computations here are done for the variables considered in the Labrador sea box as shown in Figure 2, which is different from the model dependent areas considered before. Comparison between the two regions shows no significant differences in the causal effects for the SSSMLD link in the convection mechanisms (see Fig. S3), indicating this interaction can also be identified when considering the same Labrador Sea region for all models.
@@ 4.1 Significant Links
The number of models for which each of the five links is significant are shown in Figure 4 for lags up to 10 years. Results for both the theoretical links (bottom row), as well as the influence of variables on themselves (top row), i.e. their memory, are shown. For instantaneous links also the sum of the correctly directed and unknown links is shown. Looking at the top row we see that SPG has a long memory in many models, with significant links of up to 5 years lag (lag-5) in more than 5 models. SSS, MLD and Rho have substantially shorter memory, whereas SubT-memory is two years in most models.
For the links of the proposed subpolar gyre variability mechanism (B), we find that the SPGSSS link is absent in 22 out of the 32 models (either at lag-0 or lag-1), especially when instantaneous causation is concerned (only three models, one of which also has a link at lag-1). At lag-1 eight models show a significant link. The SSSMLD link is correctly represented in 23 models, either instantaneous or at lag-1 (or both), with 13 models indicating an unknown direction for the instantaneous link. Here, 13 models show a link in the opposite direction (some with a lag-1 link in the right direction), indicating these two variables are strongly connected in models. The instantaneous MLDSubT link is also found in many models; 22 in this case with half being unoriented. At lag-1, seven models find a significant link. The SubTRho link is not present in many models, with only seven models showing a correctly directed or unoriented instantaneous link and six models indicating a significant link at lag-1. Lastly, the RhoSPG link is found in seven more models than the SPGSSS link, with eleven models (out of the 17) indicating lag-1 to be significant. In total there are six models that have both links with SPG as significant.
All but one model capture at least one step of the process (B). Only for CanESM5-1 no link is found to be significant. This is a model that shows very little convection in the Labrador sea. The best captured links are SSSMLD and MLDSubT, which are found in 23 and 27 CMIP6 models respectively. The links to and from SPG are found in only 17 and 10 models, respectively. This is to be expected as this connection relies on the modelling of advection which is dependent on the used parametrisation (Griffies, Stephen, 2018; Small et al., 2020). The SPGSSS link at longer lag is not found. There is one model, CESM2, that captures all links (except the lagged one), which will be discussed in more detail in the following section.